# TRACED: Transition-aware Regret Approximation with Co-learnability for Environment Design

### v2

@@ 3.3 Overall UED Workflow
Figure 3: Method Workflow Overview. The three panels in figure depict: (1) Task Sampling: levels are drawn from the buffer in proportion to their priority scores. (2) Priority Update: we recompute each level’s priority based on our task priority definition (Eq. 7). (3) Task Mutation: the lowest-priority levels are mutated into new variants and reinserted into the buffer.
The overall UED algorithm follows the same loop as ACCEL [7], with the sole modification that task scoring uses our task priority definition (Eq. 7) in place of positive value loss alone. We maintain a policy $\pi_{\phi}$, a level generator G, a Task Buffer $\Lambda$, and the sampling distribution $P_{D}$ over the decision variable $d\in\{0,1\}$, where $d=0$ indicates generator sampling and $d=1$ indicates buffer replay.
At each iteration $t$, we draw $d_{t}\sim P_{D}$ and proceed as follows. If $d_{t}=0$, we sample a fresh task parameter $\theta$ from the generator and run the current policy to collect a trajectory $\tau$ on $\theta$. We then compute the approximated regret score $\widehat{\mathrm{Regret}}(\tau)$ using Eq. 4. If $\widehat{\mathrm{Regret}}(\tau)$ exceeds a predefined threshold, $\theta$ is appended to $\Lambda$.
If instead $d_{t}=1$, we sample a batch of $B$ tasks $\{\theta_{k}\}_{k=1}^{B}$ from $\Lambda$ according to their TaskPriority distribution (Eq. 7). For each $\theta_{k}$, we execute the policy $\pi_{\phi}$ to collect trajectory $\tau_{k}$, perform the standard RL update on $\phi$, and compute its approximated regret $\widehat{\mathrm{Regret}}(\tau_{k})$. We then select the $N_{\text{mutate}}$ tasks with the lowest regret scores and feed each through the task editor to obtain a mutated variant $\tilde{\theta}$. Finally, we replace the corresponding entries in $\Lambda$ with these $\tilde{\theta}$, ensuring that easier tasks are systematically evolved into more challenging ones.
Following each update to the Task Difficulty Buffer, we recompute the $\mathrm{TaskDifficulty}(i,t)$ for every modified task $i$ and update the $\mathrm{CoLearnability}(j,t-1)$ for all previously sampled tasks $j$. These refreshed scores determine the priorities for the next replay phase. The algorithm thus alternates between sampling new tasks and performing prioritized replay until the policy performance converges. The full procedure is detailed in Algorithm 1.
@@ 4 Experiments
We evaluate TRACED on two procedurally generated domains: MiniGrid (MG) and BipedalWalker (BW). In both environments, we compare against DR, PLR, ACCEL (our primary baseline, since TRACED builds directly upon it), and ADD. In BW only, we additionally include the recent sota method CENIE [20]. Detailed descriptions of all baselines can be found in Appendix M.
For each domain, we track the emergent complexity of the sampled curricula during training and evaluate zero-shot test performance on held-out levels. To summarize test results, we report the median, interquartile mean (IQM), mean, and optimality gap using the rliable library [21]. The optimality gap measures how far an algorithm falls short of a target performance level, beyond which further gains are deemed negligible. Accordingly, higher IQM and lower optimality gap values indicate better performance.
All methods use Proximal Policy Optimization (PPO) [22] as the student agent. Following prior work [7, 20], we plot the performance against the number of PPO updates; the corresponding environment interactions per update are reported in Appendix C. For more details on hyperparameters and network structures, refer to Appendix F.
@@
Figure 5: Held-out Evaluation Environments. (a) Example held-out mazes for zero-shot evaluation: 4 tasks are shown (see Appendix F.1 for all 12 task definitions). (b) Example held-out BipedalWalker terrains for zero-shot evaluation: 2 tasks are shown (see Appendix F.2 for all 6 task definitions).
@@ 4.1 Partially Observable Navigation
We evaluate our curriculum design on a partially observable maze navigation domain based on MiniGrid [16], as in prior UED work. Four representatives are shown in Figure 5. The agent observes a 147-dimensional pixel observation and is trained for up to 20k PPO updates.
Figure 6: Zero-Shot Transfer Performance in MiniGrid. Aggregated solved rates on held-out MiniGrid mazes after 10k and 20k PPO updates. TRACED at 10k updates (red) matches or outperforms DR (orange), $\text{PLR}^{\perp}$ (green), ADD (gray), and ACCEL (blue) at 20k updates. Averaged over 10 random seeds.
As shown in Figure 6, our method achieves a superior solved rate at only 10k updates, matching or exceeding the 20k update performance of all baselines. In particular, TRACED’s median solved rate at 10k surpasses ACCEL at 20k, and its IQM leads the field, indicating that the majority of runs benefit rapidly from our combined regret and co-learnability scoring. The mean performance mirrors this ordering, showing that even lower-performing seeds see substantial gains. Furthermore, TRACED at 10k produces the smallest optimality gap, closing the distance to an ideal target success rate faster than any competitor; in contrast, ACCEL requires only 20k updates to halve its gap.
Table 1: Wall-clock training time comparison. Average training duration (hours $\pm$ SE over 10 runs) on the MiniGrid domain for each algorithm at 10k and 20k PPO updates.
@@
Compared to ACCEL, TRACED halves the wall-clock training time while maintaining equivalent or better transfer performance (Table 1). Even relative to ACCEL at the same 10k update budget, TRACED incurs a 6% computational overhead yet delivers a 22% relative increase in mean solved rate. Taken together, these metrics demonstrate that TRACED not only accelerates learning but also delivers more consistent and reliable zero-shot transfer across diverse MiniGrid mazes. Detailed per-task zero-shot results are provided in Appendix E, and the number of environment interactions per PPO update for each method is in Appendix C.
@@
Figure 7: Emergent maze complexity metrics. Shortest path length and number of blocks both grow faster under TRACED (pink) than ACCEL (blue). This faster ramp-up indicates that our curriculum more effectively escalates difficulty in lockstep with agent learning. Averaged over 10 random seeds with shaded 95% confidence intervals.
To analyze emergent environment complexity, we track two structural metrics for each generated maze: the length of the shortest solution path and the number of obstacles (blocks), as plotted in Figure 7. Both metrics rise much more rapidly under TRACED (red) than under ACCEL (blue), showing that our priority scoring more effectively distinguishes easy versus challenging tasks to produce a steadily escalating curriculum. Notably, the ACCEL + ATPL (brown) also drives complexity upward far faster than ACCEL. It closely parallels TRACED’s gains, demonstrating that the transition-prediction component alone contributes greatly to complexity ramp-up.
@@
Figure 8: PerfectMaze Evaluation. (a), (b) Two held-out maze instances, PerfectMazeLarge and PerfectMazeXL, used for zero-shot testing. (c) Zero-shot solved rates after 10k and 20k PPO updates (mean $\pm$ std over 10 random seeds). TRACED achieves the highest 10k performance on PerfectMazeLarge and closely matches the best 20k performance on PerfectMazeXL.
To stress-test our curriculum on extremely large, procedurally generated mazes, we introduce two PerfectMaze benchmarks. PerfectMazeLarge consists of $51\times 51$ grids with a maximum episode length exceeding 5k steps, while PerfectMazeXL scales this to $100\times 100$ grids. Figure 8 shows representative levels from each variant. We evaluate zero-shot transfer performance, measuring the mean success rate over 100 episodes per seed (10 seeds total). As depicted in Figure 8, on PerfectMazeLarge, TRACED achieves the highest 10k solved rate (27% $\pm$ 23%), outperforming ACCEL’s best 20k rate (20% $\pm$ 25%) and far exceeding ADD, DR and $\text{PLR}^{\perp}$. On the even more complex PerfectMazeXL, ACCEL narrowly leads at 12% $\pm$ 28% after 20k updates, with TRACED close behind at 10% $\pm$ 14% after just 10k updates, demonstrating that TRACED scales to extremely large mazes. Detailed per-baseline results for both benchmarks are provided in Appendix E.
@@ 4.2 Walking in Challenging Terrain
We further validate our curriculum in the continuous-control BipedalWalkerHardcore environment from OpenAI Gym [23], as modified by Wang et al. (2019) [24]. This domain features a procedurally generated terrain controlled by eight parameters, terrain roughness, pit gap frequency, stump height, stair spacing, etc., that jointly determine locomotion difficulty. We consider the complete set of eight parameters in our design space. Figure 5 illustrates two representative terrains with varying stair heights and surface roughness.
Figure 9: Aggregate zero-shot performance on BipedalWalker terrains. All scores are normalized by the maximum return of 300 in the BipedalWalker domain. TRACED at 10k updates (pink) matches or exceeds ACCEL-CENIE at 20k updates (purple) across all metrics. Averaged over 5 random seeds.
@@
Figure 10: Return progression on BipedalWalker terrains. TRACED (pink) consistently outperforms DR (orange), PLR (green), ADD (gray), and ACCEL (blue). Averaged over 10 random seeds with shaded 95% confidence intervals.
Our student agent is trained for 10k PPO updates, while baselines run up to 20k updates. We evaluate zero-shot transfer on six held-out test terrains over 100 episodes each. Figure 9 reveals consistent gains across all four aggregate metrics, median, IQM, mean, and optimality gap, on the six held-out BipedalWalker terrains. Specifically, after only 10k updates, TRACED outperforms all baselines at 20k updates. Although its variance is slightly higher, TRACED’s 10k-update curve matches or exceeds ACCEL-CENIE at 20k, demonstrating both superior sample efficiency and robust generalization. Moreover, as in MG, TRACED reaches the highest performance in half the wall-clock time on BW relative to ACCEL (see Appendix C). These results confirm that our curriculum design remains effective not only in discrete grid worlds but also in complex continuous domains. Refer to Appendix E for detailed results broken down by each terrain.
As shown in Figure 10, TRACED consistently surpasses ACCEL in zero-shot returns across all terrains. In the first 5k updates, TRACED trails ADD, DR, and PLR because it builds upon ACCEL’s seed curriculum, mutating the easiest environments first, before increasing difficulty. After 5k updates, however, TRACED continues to climb steadily, while ADD, DR, and PLR plateau in a stale regime. This sustained upward trajectory demonstrates that TRACED effectively escalates task complexity in tandem with the agent’s improving capabilities, yielding superior long-term performance.
@@ 4.3 Ablation Study
Figure 11: Both ATPL and CL are important design choices. Showing per-task zero-shot solved rates for 4 variants: full TRACED (pink), ATPL only (brown), Co-Learnability only (teal), and ACCEL (blue). Averaged over 5 random seeds.
To isolate the contributions, we compare TRACED at 10k updates against three ablations (Figure 11): ATPL only (TRACED – CL), Co-Learnability only (TRACED – ATPL), and the ACCEL baseline (TRACED – ATPL – CL). TRACED (pink) encloses the largest area in the radar chart, demonstrating that both ATPL and Co-Learnability are necessary. The ATPL only variant (brown) outperforms ACCEL (blue) on 9 out of the 12 held-out mazes, confirming that incorporating transition prediction error into the regret approximation enhances the agent’s generalization ability. In contrast, the Co-Learnability only variant (teal) falls below the ACCEL baseline, indicating that Co-Learnability is effective only when paired with our regret approximation. Detailed numerical results are provided in Appendix E. Additional ablations on BipedalWalker and sensitivity analyses for $\alpha$ and $\beta$ appear in Appendix B.
@@ 5 Conclusion
In this paper, we introduced Unsupervised Environment Design (UED) method with two key components: (i) an explicit transition prediction error term for regret approximation, and (ii) a lightweight Co-Learnability metric that captures cross-task transfer effects. By integrating these into the standard generator–replay loop, TRACED produces curricula that escalate environment complexity in tandem with agent learning.
@@
Empirically, we demonstrated on two procedurally generated domains (MiniGrid navigation and BipedalWalker) that TRACED outperforms baselines (DR, PLR, ADD, ACCEL), including sota CENIE, using only half the training updates. We showed superior zero-shot transfer success rates, faster growth in structural complexity, and scalability to extremely large mazes. Ablation studies confirmed that each component is essential: ATPL drives the primary complexity ramp-up, while co-learnability yields gains when paired with our regret estimates.
@@
Looking forward, Co-Learnability offers a simple, computationally light mechanism for capturing inter-task influences and could be further refined via more sophisticated causal estimators or learned models. More broadly, any RL setting that relies on regret-approximation stands to benefit from incorporating transition-prediction error, providing an easy-to-implement boost in sample efficiency. We anticipate that these ideas will inspire future work on adaptive curricula, advanced editing mechanisms, and broader applications of regret-guided exploration in open-ended learning environments.
Appendix A Proof
A.1 Proof of Proposition 1
Proposition 1. Assuming $\mathcal{T}_{t+1}=\Lambda$ and defining
$Y_{j}\;=\;\mathrm{TaskDifficulty}(j,t+1)\;-\;\mathrm{TaskDifficulty}(j,t),$
the co-learnability of task $i$ at time $t$ coincides with the conditional expectation, that is
$\mathrm{CoLearnability}(i,t)=\mathbb{E}_{j\sim\mathcal{T}_{t+1}}\bigl{[}Y_{j}\mid T_{i}=1,\,t\bigr{]}.$
Proof. Since $\mathcal{T}_{t+1}=\Lambda$,
$\mathbb{E}_{j\sim\mathcal{T}_{t+1}}\bigl{[}Y_{j}\mid T_{i}=1,\,t\bigr{]}\;=\;\frac{1}{|\Lambda|}\sum_{j\in\Lambda}Y_{j}=\mathrm{CoLearnability}(i,t)$
A.2 Proof of Theorem 2
Theorem 2 (Expected Staleness). Assume that at each replay step $s\geq 0$, tasks in buffer $\Lambda$ are sampled according to
$\Pr\bigl{[}i\text{ at }t_{0}+s\bigr{]}=\frac{\mathrm{TaskPriority}(i,t_{0}+s)}{\sum_{j\in\Lambda}\mathrm{TaskPriority}(j,t_{0}+s)}\,,$
[…]
Then the expected number of steps $T=t_{i}-t_{0}$ until task $i$ is sampled satisfies
$\mathbb{E}[T]\;\leq\;\frac{|\Lambda|\,M_{0}}{\bigl{(}\mathrm{TaskPriority}(i,t_{0})\bigr{)}}\,.$
Proof. Let
$q_{s}\;=\;\Pr\bigl{[}\text{choose }i\text{ at time }t_{0}+s\bigr{]}\;=\;\frac{\mathrm{TaskPriority}(i,t_{0}+s)}{\sum_{j\in\Lambda}\mathrm{TaskPriority}(j,t_{0}+s)}\,,$
so that $1-q_{s}$ is the probability $i$ is not chosen at step $s$.
Since for every $j\in\Lambda$ we have $\mathrm{TaskPriority}(j,t_{0}+s)\leq M_{0})$, it follows that
$\sum_{j\in\Lambda}\mathrm{TaskPriority}(j,t_{0}+s)\;\leq\;|\Lambda|\;M_{0}.$
Hence
$q_{s}\;=\;\frac{\mathrm{TaskPriority}(i,t_{0}+s)}{\sum_{j}\mathrm{TaskPriority}(j,t_{0}+s)}\;\geq\;\frac{\mathrm{TaskPriority}(i,t_{0}+s)}{|\Lambda|\;M_{0}}\;\geq\;\frac{\mathrm{TaskPriority}(i,t_{0})}{|\Lambda|\;M_{0}}=:\;c,$ (8)
For any nonnegative integer $k$,
$\Pr[T>k]$ $=\Pr\bigl{(}\text{task }i\text{ is not chosen in any of steps }0,1,\dots,k-1\bigr{)}$
$=\prod_{s=0}^{k-1}\bigl{(}1-q_{s}\bigr{)}\;\leq\;\prod_{s=0}^{k-1}\bigl{(}1-c\bigr{)}=\bigl{(}1-c\bigr{)}^{k}.$
where the inequality uses (8).
$\mathbb{E}[T]=\sum_{k=0}^{\infty}\Pr[T>k]\;\leq\;\sum_{k=0}^{\infty}(1-c)^{k}\;=\;\frac{1}{c}\;=\;\frac{|\Lambda|\;M_{0}}{\bigl{(}\mathrm{TaskPriority}(i,t_{0})\bigr{)}}.$
This completes the proof.
Appendix B Additional Experiments
B.1 Ablation Study of Scaling Factor
In this ablation, we keep every setting in Appendix H fixed except for one of the weight-scaling factors, $\alpha$ (ATPL weight), $\beta$ (Co-Learnability weight). Deviating from defaults on either $\alpha$ or $\beta$ reduces the solved rates in most domains as shown in Table 2. For both ignoring one of the weighted terms ($\alpha=0.0$ or $\beta=0.0$) and using excessively large weight ($\alpha=100.0$ or $\beta=100.0$), solved rates show slight improvement for some individual tasks, but statistical drops for overall performance. The default TRACED weights show the best performance compared to fixed projection baselines. This result shows that an effective balance between ATPL and Co-Learnability is needed.
Table 2: MiniGrid Ablation Solved Rates. Mean $\pm$ SE of solved rates on 12 held-out MiniGrid tasks, comparing different fixed projection Weight baselines ($\alpha=0.0$, $\alpha=100.0$, $\beta=0.0$, $\beta=100.0$) and TRACED ($\alpha=1.0$, $\beta=1.0$) over 5 seeds. For TRACED, 5 runs were randomly sampled from 10 seeds to ensure consistent evaluation. Each method is evaluated after 10k PPO updates. Bold indicates the best result per task; underline indicates the second-best.

### v3

@@ 3.3 Overall UED Workflow
Figure 3: Method Workflow Overview. The three panels in figure depict: (1) Task Sampling: levels are drawn from the buffer based on their priority scores. (2) Priority Update: we recompute each level’s priority based on our task priority definition (Eq. 7). (3) Task Mutation: the lowest‑priority levels are mutated into new variants and reinserted into the buffer.
The overall UED algorithm follows the ACCEL loop (Parker-Holder et al., 2022), with the sole change that task scoring uses our Task Priority (Eq. 7) in place of Positive Value Loss (PVL) alone (Figure 3). At each curriculum update time $t$, with probability $d$ we sample a new level uniformly at random, and with probability 1d1-d we sample a level from the replay buffer according to its Task Priority score. Early in training, when the buffer is empty, we sample levels uniformly at random for a warm-up period. The warm-up length matches ACCEL’s setting.
After training the agent on the selected level, we approximate its regret using the agent’s value-related loss and the transition-prediction loss (for Task Difficulty), and append the result to the task difficulty buffer. Co-Learnability is updated following Eq. 6 after the next curriculum step, once the newly visited levels have had their difficulties updated. The updated Task Difficulty and Co-Learnability define Task Priority via Eq. 7, which then governs both new-task sampling and prioritized replay at future times. The algorithm alternates between sampling new tasks and prioritized replay until policy performance converges. The full procedure is given in Algorithm 1.
@@ 4 Experiments
We evaluate TRACED on two procedurally generated domains: MiniGrid (MG) and BipedalWalker (BW). In both environments, we compare against DR, PLR, ACCEL (our primary baseline, since TRACED builds directly upon it), and ADD. In BW only, we additionally include the recent sota method CENIE (Teoh et al., 2024). We exclude SFL (Rutherford et al., 2024) because, on MiniGrid, it does not outperform PLR (Rutherford et al., 2024). Detailed descriptions of all baselines can be found in Appendix K.
For each domain, we track the emergent complexity of the sampled curricula during training and evaluate zero‐shot test performance on held‐out levels. To summarize test results, we report the median, interquartile mean (IQM), mean, and optimality gap using the rliable library (Agarwal et al., 2021). The optimality gap measures how far an algorithm falls short of a target performance level, beyond which further gains are deemed negligible. Accordingly, higher IQM and lower optimality gap values indicate better performance.
All methods use Proximal Policy Optimization (PPO) (Schulman et al., 2017) as the student agent. Following prior work (Parker-Holder et al., 2022; Teoh et al., 2024), we plot performance versus the number of PPO updates and report the corresponding environment interactions per update in Appendix F. Our only deviation from ACCEL is the number of PPO workers (i.e., parallel trajectory collectors), chosen for computational constraints: MiniGrid uses 16 workers instead of 32, and BipedalWalker uses 4 instead of 16. Consequently, some baseline scores may differ from those in the ACCEL paper. However, Appendix A.4 shows that TRACED has no worker-specific tuning and that reducing workers does not render the comparison unfair. We report TRACED at 10k updates (rather than 20k) because single-seed, long-horizon runs (45k PPO updates) on MiniGrid reveal post-convergence oscillations in both TRACED and ACCEL, and reporting earlier avoids this confound (Appendix B). Additional hyperparameters and architectures are provided in Appendix H.
@@
Figure 5: Held‑out Evaluation Environments. (a) Example held‑out MiniGrid mazes for zero‑shot evaluation: 4 tasks are shown (see Appendix H.1 for all 12 task definitions). (b) Example held‑out BipedalWalker terrains for zero‑shot evaluation: 2 tasks are shown (see Appendix H.2 for all 6 task definitions).
@@ 4.1 Partially Observable Navigation
We evaluate our curriculum design on a partially observable maze navigation domain based on MiniGrid (Chevalier-Boisvert et al., 2023), as in prior UED work. Four representatives are shown in Figure 5. The agent observes a 147‑dimensional pixel observation and is trained for up to 20k PPO updates.
Figure 6: Zero‑Shot Transfer Performance in MiniGrid. Aggregated solved rates on held‑out MiniGrid mazes after 10k and 20k PPO updates. TRACED at 10k updates outperforms baselines at 20k updates.
Our method achieves a superior solved rate at only 10k updates, matching or exceeding the 20k update performance of all baselines (Figure 6). In particular, TRACED’s median solved rate at 10k surpasses ACCEL at 20k, and its IQM leads the field, indicating that the majority of runs benefit rapidly from our combined regret and Co‑Learnability scoring.
Table 1: Wall-clock training time comparison. Average training duration (hours $\pm$ SE over 10 runs) on the MiniGrid domain.
@@
Compared to ACCEL, TRACED halves the wall‑clock training time while maintaining equivalent or better transfer performance (Table 1). Even relative to ACCEL at the same 10k update budget, TRACED incurs a 6% computational overhead yet delivers a 22% relative increase in mean solved rate. Taken together, these metrics demonstrate that TRACED not only accelerates learning but also delivers more consistent and reliable zero‑shot transfer across diverse MiniGrid mazes. Detailed per‑task zero‑shot results are provided in Appendix P, and the number of environment interactions per PPO update for each method is in Appendix F.
@@
Figure 7: Emergent maze complexity metrics. Shortest path length and number of blocks both grow faster under TRACED (pink) than ACCEL (blue). This faster ramp‑up indicates that our curriculum more effectively escalates difficulty in lockstep with agent learning.
To analyze emergent environment complexity, we track two structural metrics for each generated maze: (i) the length of the shortest solution path and (ii) the number of obstacles (Figure 7). Curves are averaged over 10 seeds with shaded 95% confidence intervals. Both metrics increase substantially faster under TRACED than under ACCEL, indicating that our priority scoring more effectively separates easy from challenging tasks and yields a steadily escalating curriculum. Notably, ACCEL+ATPL also drives complexity upward far faster than ACCEL alone and closely tracks TRACED, demonstrating that the transition-prediction component on its own contributes strongly to complexity ramp-up.
@@
Figure 8: PerfectMaze Evaluation. (a), (b) Two held‑out maze instances, PerfectMazeLarge and PerfectMazeXL, used for zero‑shot testing. (c) Zero‑shot solved rates. TRACED achieves the highest 10k performance on PerfectMazeLarge and closely matches the best 20k performance on PerfectMazeXL.
To stress‐test our curriculum on extremely large, procedurally generated mazes, we introduce two PerfectMaze benchmarks. PerfectMazeLarge consists of $51\times 51$ grids with a maximum episode length exceeding 5k steps, while PerfectMazeXL scales this to $100\times 100$ grids. Figure 8 shows representative levels from each variant. We evaluate zero‐shot transfer performance, measuring the mean success rate over 100 episodes per seed (10 seeds total).
On PerfectMazeLarge, TRACED achieves the highest 10k solved rate (27% $\pm$ 23%), outperforming ACCEL’s best 20k rate (20% $\pm$ 25%) and far exceeding ADD, DR and $\text{PLR}^{\perp}$ (Figure 8). On the even more complex PerfectMazeXL, ACCEL narrowly leads at 12% $\pm$ 28% after 20k updates, with TRACED close behind at 10% $\pm$ 14% after just 10k updates, demonstrating that TRACED scales to extremely large mazes. Detailed per-baseline results for both benchmarks are provided in Appendix P.
@@ 4.2 Walking in Challenging Terrain
We further validate our curriculum in the continuous‐control BipedalWalkerHardcore environment from OpenAI Gym (Brockman et al., 2016), as modified by Wang et al. (2019). This domain features a procedurally generated terrain controlled by eight parameters, terrain roughness, pit gap frequency, stump height, stair spacing, etc., that jointly determine locomotion difficulty. We consider the complete set of eight parameters in our design space. Figure 5 illustrates two representative terrains with varying stair heights and surface roughness. We evaluate zero‑shot transfer on six held‑out test terrains over 100 episodes each, averaged over five random seeds.
Figure 9: Aggregate zero‑shot performance on BipedalWalker terrains. All scores are normalized by the maximum return of 300 in the BipedalWalker domain. TRACED at 10k updates (pink) matches or exceeds ACCEL-CENIE at 20k updates (purple) across all metrics.
@@
Figure 10: Return progression on BipedalWalker terrains. TRACED consistently outperforms baselines.
TRACED delivers consistent gains across all four aggregate metrics, median, interquartile mean (IQM), mean, and optimality gap, on the six held-out BipedalWalker terrains (Figure 9). After only 10k updates, it already outperforms all baselines evaluated at 20k. As in MiniGrid, TRACED also reaches its peak performance in roughly half the wall-clock time of ACCEL on BipedalWalker (Appendix F). Moreover, TRACED consistently surpasses ACCEL in zero-shot returns across all terrains throughout training (Figure 10), further underscoring the effectiveness of its curriculum design. Detailed numerical results are provided in Appendix P.
@@ 4.3 Ablation Study
Figure 11: Ablation Study on TRACED. Both ATPL and CL are important design choices.
To isolate component contributions, we compare TRACED at 10k updates with two ablations (Figure 11): ATPL only (TRACED - CL) and Co-Learnability only (TRACED - ATPL), each evaluated with five seeds. Results are reported following Agarwal et al. (2021). On MiniGrid, TRACED outperforms both variants across four metrics, indicating that each component contributes meaningfully to performance. The same trend holds on BipedalWalker (Appendix A). Ablations on the scaling factors $\alpha$ and $\beta$ further show that both elements play significant roles in TRACED and that carefully balancing ATPL and CL can yield additional gains (Appendix A.1). Detailed numerical results are provided in Appendix P.
4.4 Analysis on Curriculum Progression
Table 2: Proportions (%) of Easy, Moderate, and Challenging levels in the level buffer during PPO updates.
Method Difficulty 0k 5k 10k 15k 20k
ACCEL Easy $100$ $100$ $79.2$ $26.7$ $9.2$
Moderate 0 0 $20.8$ $73.3$ $90.8$
Challenging 0 0 0 0 0
TRACED Easy $100$ $72$ $24.5$ $19.5$ $11.2$
Moderate 0 $25.7$ $60.9$ $68$ $77.3$
Challenging 0 $2.3$ $14.6$ $12.5$ $11.5$
To analyze how the curriculum evolves, we examine the BipedalWalker level buffer over time and categorize each generated level into three difficulty bands: Easy, Moderate, and Challenging, based on environment hyperparameters. Following Teoh et al. (2024), we set thresholds for Stump Height (2.4), Pit Gap (6), Ground Roughness (4.5), and Stairs Height (5). Levels exceeding no thresholds are labeled Easy, those exceeding exactly one threshold are Moderate, and those exceeding at least two thresholds are Challenging. Table 2 reports the proportion of each difficulty in the buffer at different PPO update steps.
TRACED progressively shifts mass from Easy to Moderate and Challenging, introducing nontrivial proportions of Challenging levels by 10k (14.6%) and maintaining them thereafter (11.5% at 20k). In contrast, ACCEL never surfaces Challenging levels even at 20k (0% throughout), with the buffer dominated by Moderate levels by 20k (90.8%). This steady escalation under TRACED aligns with its faster convergence and stronger performance already at 10k updates. The respective contributions of ATPL and Co-Learnability (CL) to curriculum progression are analyzed in Appendix C, and a visualization of level evolution is provided in Appendix D.
@@ 5 Conclusion
In this paper, we introduced Unsupervised Environment Design (UED) method with two key components: (i) an explicit transition-prediction error term for regret approximation, and (ii) a lightweight Co‑Learnability metric that captures cross‐task transfer effects. By integrating these into the standard generator-replay loop, TRACED produces curricula that escalate environment complexity in tandem with agent learning.
@@
Empirically, we demonstrated on two procedurally generated domains (MiniGrid navigation and BipedalWalker) that TRACED outperforms baselines (DR, PLR, ADD, ACCEL), including sota CENIE, using only half the training updates. We showed superior zero‑shot transfer success rates, faster growth in structural complexity, and scalability to extremely large mazes. Ablation studies confirmed that each component is essential: ATPL drives the primary complexity ramp‑up, while Co‑Learnability yields gains when paired with our regret estimates.
@@
Looking forward, Co‑Learnability offers a simple, computationally light mechanism for capturing inter‐task influences and could be further refined via more sophisticated causal estimators or learned models. More broadly, any RL setting that relies on regret‑approximation stands to benefit from incorporating transition‐prediction error, providing an easy‐to‐implement boost in sample efficiency. We anticipate that these ideas will inspire future work on adaptive curricula, advanced editing mechanisms, and broader applications of regret‐guided exploration in open‑ended learning environments.
Appendix A Additional Experiments
A.1 Ablation Study on the Scaling Factor
In this ablation, we keep every setting in Appendix O fixed except for one of the weight-scaling factors, $\alpha$ (ATPL weight), $\beta$ (Co-Learnability weight). Results are averaged over five random seeds and reported as mean $\pm$ standard error on 12 held-out MiniGrid tasks. Each method is evaluated after 10k PPO updates.
Table 3: Ablation study on the scaling factor. Comparing different fixed projection Weight baselines ($\alpha=0.0$, $\alpha=100.0$, $\beta=0.0$, $\beta=100.0$) and TRACED ($\alpha=1.0$, $\beta=1.0$). Bold indicates the best; underline indicates the second-best.